OPTICAL SELECTION OF SLOW MOLECULES USING TRANSIENT LINE NARROWING, APPLICA-TION TO SF_{6}

\underline{W. VON KLITZING}, Laboratoire Kastler Brossel, Ecole Normale Supérieure, 75231 Paris Cedex 05, France; R.J. BUTCHER, Cavendish Laboratory, University of Cambridge, Cambridge CB3 0HE, UK.

A novel effect has been found which reduces the linewidth of absorptions to below their linewidth defined by the limited average interaction time between the molecules and the laser light. For transit-time broadened transitions it will be demonstrated that transient line narrowing can reduce the linewidth by a factor of five whilst increasing signal to noise ratio. Using the Cambridge 10μm CO_{2}-Laser Spectrometer absorptions in room temperature SF_{6} have been recorded with a half-width-half-maximum of 1.4 kHz which implies a transverse velocity distribution with an effective temperature of only 1K. This method practically eliminates shifting and broadening effects like the second order Doppler effect.